# Recursive Inference Scaling: A Winning Path to Scalable Inference in Language and Multimodal Systems

### v2

@@ Abstract
Recent research in language modeling reveals two scaling effects: the well-known improvement from increased training compute, and a lesser-known boost from applying more sophisticated or computationally intensive inference methods. Inspired by recent findings on the fractal geometry of language, we introduce Recursive INference Scaling (RINS) as a complementary, plug-in recipe for scaling inference time. For a given fixed model architecture and training compute budget, RINS substantially improves language modeling performance. It also generalizes beyond pure language tasks, delivering gains in multimodal systems, including a +2% improvement in 0-shot ImageNet accuracy for SigLIP-B/16. Additionally, by deriving data scaling laws, we show that RINS improves both the asymptotic performance limits and the scaling exponents. These advantages are maintained even when compared to state-of-the-art recursive techniques like the “repeat-all-over” (RAO) strategy in Mobile LLM. Finally, stochastic RINS not only can enhance performance further but also provides the flexibility to optionally forgo increased inference computation at test time with minimal performance degradation.
Machine Learning, ICML
@@ 1 Introduction
There has been a proliferation of research in recent years pointing to the pivotal role of scale, and how its benefits could be predicted empirically (Mukherjee et al., 2003; Hestness et al., 2017; Johnson et al., 2018; Rosenfeld et al., 2019; Kaplan et al., 2020; Ghorbani et al., 2021; Alabdulmohsin et al., 2022; Bansal et al., 2022; Zhai et al., 2022). Generally, the performance of deep neural networks $f(x)$ (such as its error rate or log-perplexity) often follows a power law $f(x)\sim\beta x^{-c}+\varepsilon$ as one varies a dimension of interest $x$, such as the size of the training data or model parameters. These “scaling laws,” as they are known in the literature, have been used, among others, to determine the training data size needed to achieve a specified level of accuracy (Cho et al., 2015; Beleites et al., 2013; Figueroa et al., 2012) and to optimize the model architecture (Kaplan et al., 2020; Hoffmann et al., 2022; Alabdulmohsin et al., 2024b). They have also been justified theoretically (Bahri et al., 2021; Hutter, 2021; Sharma & Kaplan, 2022).
@@
Figure 1: left: Illustration of Recursive INference Scaling (RINS) compared to a standard forward pass. In RINS, the model $f:\mathcal{X}\to\mathcal{Y}$ is split into two parts: the first block $f_{A}:\mathcal{X}\to\mathcal{X}$ is applied iteratively to its own output for $r$ iterations before passing the output to the second block. right: Illustrative examples of models with different signatures and degrees. From top to bottom: (1) Baseline (Signature: AB, Degree: 1), a standard feedforward architecture with no recursion. (2) repeat-all-over (RA) (Liu et al., 2024), where the entire model is recursively applied on its output. When recursion is done twice, it has a signature of ABAB. (3) RINS with signature A3B. (4) (A3B)2 whose degree is 2, in which the same parameter sharing signature is applied on each of the two blocks A and B.
@@
Besides this conventional approach of scaling the training compute, the impact of increased inference compute on model performance has emerged as another key scaling dimension. Whereas OpenAI’s GPT-o3 has prominently demonstrated this recently (OpenAI, 2024), earlier work had already explored this direction. For example, chain-of-thought (CoT) prompting has shown that eliciting longer inference paths through additional token generation could improve the reasoning capabilities of language models (Wei et al., 2024). This is similar to the demonstrated success of critiquing before evaluating (Ankner et al., 2024). Also, AlphaCode (Li et al., 2022) and Codex (Chen et al., 2021) generate multiple samples during inference in order to enhance code generation performance. Remarkably,  Brown et al. (2024) shows that the benefit of sampling multiple solutions in tasks such as mathematics and coding—when measured by coverage—holds for up to four orders of magnitude of inference compute. These approaches suggest that, like training compute, inference compute follows systematic scaling patterns that can be leveraged to improve models. Refer to the survey (Welleck et al., 2024) for more details.
@@
Recently, it has also been noted that language exhibits a “self-similar” (fractal) nature, meaning that similar patterns repeat across different scales of its representation, from individual words to entire sentences and paragraphs (Alabdulmohsin et al., 2024a). Inspired by this finding, we examine if model recursion, which can be interpreted as a particular form of scale-invariant decoding, offers a complementary approach for scaling inference time in language models. To this end, we examine an extensive set of parameter-sharing strategies and, indeed, identify the best to be a special form of recursion, which we term Recursive INference Scaling (RINS). We show that RINS yields significant performance gains compared to other parameter-sharing methods, even when controlling for model size and training compute.
@@
RINS builds upon the concept of model recursion but recasts it as a powerful inference-time scaling strategy. It leverages a simple yet profound idea: use your existing architecture and training compute budget as is, but exploit the self-similar structure of language by recursively applying an early portion of your network to refine its output. In turn, this simple strategy improves performance significantly, with the primary cost being a scaled inference time.
Recursion has shown promise in language modeling, with recent work by Liu et al. (2024) demonstrating that recursive architectures outperform similarly sized vanilla models trained on the same number of tokens. Notably, Liu et al. (2024) finds that a simple “repeat-all-over” (RAO) approach, where the entire model is recursively applied, surpasses more complex parameter-sharing strategies in language.
@@
However, while Liu et al. (2024) demonstrate the sample efficiency of recursive architectures, in which models are compared when trained on the same number of tokens, their analysis does not explicitly account for the increased computational cost of recursive operations during training. Hence, it remains unclear whether the performance gains observed in prior work come from the inherent advantages of model recursion or simply from having increased the training compute. Indeed, our findings suggest that for moderately sized models (over 1 billion parameters), the performance gains of the “repeat-all-over” (RAO) strategy can be matched by extending the training of a standard model to consume an equivalent compute. RINS, by contrast, significantly outperforms all other baselines on a compute- and parameter-matched setup, including when scaling inference by increasing the context length (see Figure 2).
@@
Crucially, a stochastic variant of RINS, in which the number of recursion rounds is sampled at training time from a binomial distribution, not only can enhance performance further, such as in multimodal systems, but also provides the flexibility to optionally forgo increased inference computation at test time with minimal performance degradation. See Section 4 for details.
Efficiency.
@@
We conduct our experiments mostly on small models, which are typically intended for deployment environments with stringent memory limitations (Liu et al., 2024). Given the direct relation between a model’s memory footprint and its parameter count (e.g. a 1 billion parameter model with 16-bit floating-point precision requires 2GB of DRAM), the ability to enhance accuracy while maintaining a fixed parameter count is highly desirable. RINS achieves this by enabling significant performance gains without increasing parameter count for the same training compute.
@@
    propose a taxonomy of parameter-sharing architectures, empirically evaluating their effectiveness across multiple domains. Our comprehensive analysis identifies RINS as a powerful approach, outperforming methods like RAO used in Mobile LLM (Liu et al., 2024), and scaling inference by increasing the sequence length.
@@
    argue that RINS’ performance gain likely stems from the self-similar (fractal) nature of language, as a similar analysis in vision yields minimal improvements.
@@
    use RINS to extend inference scaling beyond language to multimodal systems that incorporate language in their processing, such as contrastive models. In particular, our SigLIP-RINS-B/16 outperforms the popular SigLIP-B/16 (Zhai et al., 2023) by a wide margin; e.g. improving 0-shot accuracy in ImageNet from 77.3% to 79.6% and CIFAR100 from 70.3% to 80.7%.
@@
  5. 5.
    show that stochastic RINS can enhance performance even further, such as in multimodal systems, while offering the option to revert to non-recursive inference at test time with minimal performance degradation.
@@ 2 Recursive Inference Scaling
2.1 Overview
Before describing how RINS works, we briefly formalize definitions. Let $\mathcal{X}$ be a fixed domain, often the space of sequences of soft tokens of embedding dimension $d$. Let $\mathbb{L}=\{l_{1},l_{2},...,l_{n}\}$ be a fixed set of $n$ unique blocks, where each block $l_{k}:\mathcal{X}\to\mathcal{X}$ is a function mapping from the input space $\mathcal{X}$ to the same output space $\mathcal{X}$. By “unique” here we simply imply that such blocks (which typically comprise of multiple layers each) are not constrained to share the same parameters. Let $\mathcal{G}(\mathbb{L})$ be the space of all possible computation graphs representing neural network architectures that can be constructed by composing blocks from the set $\mathbb{L}$, while $f\in\mathcal{G}(\mathbb{L})$ be one specific architecture.
Figure 1 (right) illustrates some examples for the case when $|\mathbb{L}|=2$. For instance, one can repeatedly apply the entire model, as in the “repeat-all-over” (RAO) approach in Mobile LLM (Liu et al., 2024), or recursively apply a strict subset of the architecture, such as a single a block within the model. The choice of arrangement of blocks can significantly impact the model’s performance and efficiency.
Formally, let $C(f)$ be the actual computational cost (in FLOPs) of training $f\in\mathcal{G}(\mathbb{L})$ on a dataset sampled i.i.d. from distribution $\mathcal{D}$, considering only the forward pass. Also, $\mathcal{L}(f)$ is a performance metric of interest (e.g., validation loss or error rate) for model $f$. We assume that a lower value of $\mathcal{L}$ indicates better performance.
@@ 2.1 Overview
For a fixed set of unique blocks $\mathbb{L}$ and a training compute budget $c$, a recursive architecture $f^{\star}\in\mathcal{G}(\mathbb{L})$ is considered “better” than another architecture $f\in\mathcal{G}(\mathbb{L})$ if $C(f^{\star})\leq C(f)\leq c$ and $\mathbb{E}[\mathcal{L}(f^{\star})]\leq\mathbb{E}[\mathcal{L}(f)]$.
@@
Model recursion offers a simple, plug-in approach for scaling inference time. By applying some layers iteratively to their own output, we effectively increase the computational path length during inference without altering the underlying model architecture. This allows us to exploit the benefits of increased inference compute. Importantly, it is complementary to other techniques like chain-of-thought (CoT) prompting (Wei et al., 2024), which achieve performance gains by eliciting longer generation sequences, and repeated sampling (Li et al., 2022; Chen et al., 2021).
Figure 2: Language models are trained on 200B tokens. The $x$-axis is the training cost in units of layer $\times$ step. Notably, the performance advantage of RINS increases with increased computational budget. The long-sequence baseline, using a context length of 1,536 tokens, exhibits lower performance due to processing fewer examples to maintain the same FLOPs count. See Figure 4 for results with extended training duration and Figure 5 for results with larger (1B parameter) models, further demonstrating the benefits of RINS.
@@
As discussed in Section 1, language has been shown to exhibit self-similarity, meaning that similar patterns repeat across multiple levels of granularity, from the structure of individual sentences to the composition of entire texts (Alabdulmohsin et al., 2024a). This observation suggests that recursive (scale-invariant) decoding could be particularly beneficial for language processing.
To determine the extent to which this is true, we systematically explore a vast space of parameter sharing strategies. We use the following taxonomy based on two key concepts: (1) signature and (2) degree. The “signature” of an architecture describes how parameter sharing is applied across different blocks of the model. For example, an architecture with the signature AB2C indicates that the model is partitioned into three unique blocks (A, B, and C) of equal size. The processing flow in this architecture would be: apply block A on the input, apply block B to A’s output, apply block B again to its own output (hence the exponent), and finally apply C to the output of B. Note that while the signature AB2C has the same parameter count as a simpler ABC architecture, it incurs approximately 33% more compute due to the repeated application of block B. This highlights the trade-off between computational cost and performance using recursion that we will carefully study in this work.
On the other hand, “degree” specifies whether recursion is applied only at the level of entire blocks, or if it extends to individual components within blocks, as illustrated in Figure 1 (bottom). A degree of 2, for example, means that each of the blocks A, B, and C in our example have themselves the same recursive pattern used in the overall model. This adds another dimension to our taxonomy, allowing for a finer-grained classification of recursive architectures. Note that degree makes notation concise, but is not necessary; e.g. (ABB)2 is equivalent to ABBCDDCDD. By systematically varying both signature and degree, we can comprehensively explore the space of recursive models and identify optimal configurations. Figure 3 provides a detailed pseudocode.
With this framework, we now state our main investigation:
Which family of architectures (i.e. signatures and degrees) lead to better performance according to Definition 2.1 under fixed compute budget?
To reiterate, this is a non-trivial question because it is conceivable for a non-recursive model to outperform all others given that it sees more training examples, utilizing the compute budget without resorting to recursion. For instance, increasing the context length in language modeling can be inferior to training on more examples within a limited compute budget. We present examples of this in Section 3.
Nevertheless, our analysis reveals that Recursive INference Scaling (RINS) emerges as a clear winner. Its consistent superiority suggests that it captures a fundamental principle for efficiently processing language. We hypothesize that this is due to the self-similar geometry of language.
@@
RINS corresponds to architectures with degree 1 and signature ArB, for some integer $r>1$.
In other words, RINS partitions a given model into two equally-sized blocks A and B. For example, A would correspond to the first 6 layers while B would correspond to the second 6 layers in a 12-layer neural network (ignoring the embedding layer and classifier head). The first block A is recursively applied on its own output $r$ times before passing the output to B. See Figure 1 for illustration and Figure 3 for a detailed pseudocode. In Section 6, we show how the optimal value of $r$ depends on the allocated compute budget.
@@
Our claim can be summarized as follows. Once a language model is chosen and the training compute budget is planned, one should enable RINS, which does not change the model size, and train the new RINS-enabled model for the same amount of compute. This is achieved by partitioning the original model into two equally-sized blocks and applying recursion to the first block alone. Thus, neither model size nor training compute is affected. Our empirical results demonstrate that RINS-enabled models will consistently outperform baseline models. In addition, stochastic RINS with some positive skip probability $p_{s}>0$ (see Section 4) offers the option of forgoing scaling inference at test time with little performance degradation.

### v3

@@ Abstract
Inspired by recent findings on the fractal geometry of language, we introduce Recursive INference Scaling (RINS) as a complementary, plug-in recipe for scaling inference time in language and multimodal systems. RINS is a particular form of recursive depth that significantly outperforms +55 other variants, including the recent “repeat-all-over” (RAO) strategy in Mobile LLM (Liu et al.,, 2024) and latent recurrent thinking (Geiping et al.,, 2025). Unlike prior works, we carry out our comparisons on a compute-matched regime, and demonstrate that for a fixed model size and training compute budget, RINS substantially improves language modeling performance. It also generalizes beyond pure language tasks, delivering gains in multimodal systems, including a +2% improvement in 0-shot ImageNet accuracy for SigLIP-B/16. Additionally, by deriving data scaling laws, we show that RINS improves both the asymptotic performance limits and the scaling exponents. More importantly, with light-weight (linear) adapters (comprising $<1\%$ of model parameters) and stochastic dropout, RINS offers a no-regret strategy, meaning that RINS-enabled pretraining improves performance in language modeling even when recursive depth is not applied at inference time. This corresponds to improving performance on a training compute-, parameter-, and inference-matched regime, suggesting its potential as a viable component of LLM pretraining!
@@ 1 Introduction
There has been a proliferation of research in recent years pointing to the pivotal role of scale, and how its benefits could be predicted empirically (Hestness et al.,, 2017; Kaplan et al.,, 2020; Alabdulmohsin et al.,, 2022; Bansal et al.,, 2022; Zhai et al.,, 2022). Generally, the performance of deep neural networks $f(x)$ (such as its error rate or log-perplexity) often follows a power law $f(x)\sim\beta x^{-c}+\varepsilon$ as one varies a dimension $x$, such as the data size or model parameters. These “scaling laws,” as they are known today, have been used, among others, to determine the training data size needed for a specified level of accuracy (Cho et al.,, 2015; Beleites et al.,, 2013; Figueroa et al.,, 2012) and to optimize the model architecture (Kaplan et al.,, 2020; Hoffmann et al.,, 2022; Alabdulmohsin et al., 2024b, ), with some theoretical justification (Bahri et al.,, 2021; Hutter,, 2021; Sharma and Kaplan,, 2022).
@@
Figure 1: left: In RINS, the model $f:\mathcal{X}\to\mathcal{Y}$ is split into two parts: the first block $f_{A}:\mathcal{X}\to\mathcal{X}$ is applied iteratively to its own output $r$ times before passing the output to the second block. right: Illustrative examples of models with different signatures and degrees. From top to bottom: (1) Baseline (Signature: AB, Degree: 1), a feedforward architecture with no recursion. (2) repeat-all-over (RA) (Liu et al.,, 2024), where the entire model is recursively applied on its output. When recursion is done twice, it has a signature of ABAB. (3) RINS with signature A3B. (4) (A3B)2 whose degree is 2, in which the same parameter sharing signature is applied on each of the two blocks A and B.
@@
Besides this conventional approach of scaling training compute, the impact of increased inference compute on model performance has emerged as another key scaling dimension. For example, chain-of-thought (CoT) prompting show that eliciting longer inference paths through additional token generation could improve the reasoning capabilities of LLMs (Wei et al.,, 2024), similar to the success of critiquing before evaluating (Ankner et al.,, 2024). Also, AlphaCode (Li et al.,, 2022) and Codex (Chen et al.,, 2021) generate multiple samples during inference to enhance code generation. Remarkably,  Brown et al., (2024) shows that the benefit of sampling multiple solutions in tasks such as mathematics and coding—when measured by coverage—holds for up to four orders of magnitude of inference compute. Thus, inference compute follows systematic scaling patterns that can be leveraged to improve models. Refer to the survey by Welleck et al., (2024) for more details.
@@
Recently, it has also been noted that language exhibits a “self-similar” (fractal) nature, meaning that similar patterns repeat across different scales of its representation, from individual words to entire sentences and paragraphs (Alabdulmohsin et al., 2024a, ). Inspired by this finding, we examine if recursive depth, which can be interpreted as a form of scale-invariant decoding, offers a complementary approach for scaling inference time in language models. To this end, we examine an extensive set of parameter-sharing strategies and, indeed, identify the best to be a special form of recursion, which we term Recursive INference Scaling (RINS). We show that RINS yields significant performance gains over +55 other methods when controlling for model size and training compute.
@@
RINS builds upon the concept of model recursion but recasts it as a powerful inference-time scaling strategy. It leverages a simple yet profound idea: use your existing architecture and training compute budget as is, but exploit the self-similar structure of language by recursively applying an early portion of your network to refine its output. In turn, this simple strategy improves performance significantly.
@@
Recursion has shown promise in language modeling, with recent work by Liu et al., (2024) and Geiping et al., (2025) demonstrating that recursive architectures outperform similarly sized vanilla models trained on a similar number of tokens. However, while such works demonstrate the sample efficiency of recursive architectures, in which models are compared when trained on a similar number of tokens, their analysis does not explicitly account for the increased computational cost of recursive operations during training. Hence, it remains unclear whether the performance gains observed in prior work come from the inherent advantages of model recursion or simply from having increased the training compute. Indeed, our findings suggest that for moderately sized models (over 1 billion parameters), the performance gains of “repeat-all-over” (RAO) in MobileLLM (Liu et al.,, 2024) can be matched by training the baseline model longer to consume an equivalent compute. RINS, by contrast, significantly outperforms all other baselines on a compute- and parameter-matched setup, including when scaling inference by increasing the context length (see Figure 2).
@@
Crucially, a stochastic variant of RINS not only can enhance performance further, such as in multimodal systems, but also provides the flexibility to optionally forgo increased inference computation at test time with minimal performance degradation. We show that combining stochastic RINS with lightweight (linear adapters), comprising $<1\%$ of parameters, offers a no-regret strategy, meaning that RINS-enabled pretraining improves performance in langauge modeling even when recursive depth is not applied at inference time! See Section 4 for details.
@@
We conduct our experiments mostly on compact models, which are typically intended for deployment environments with stringent memory limitations (Liu et al.,, 2024). Given the direct relation between a model’s memory footprint and its parameter count (e.g. a 1 billion parameter model with 16-bit floating-point precision requires 2GB of DRAM), the ability to enhance accuracy while maintaining a fixed parameter count is highly desirable. RINS achieves this by unlocking significant gains without increasing parameter count for the same training compute. In Section 6, we study the effect of sharing the KV cache during recursion to reduce memory footprint even further. While KV cache sharing diminishes some of the gain, RINS with KV cache sharing still enjoys an advantage.
@@
    propose a taxonomy of parameter-sharing architectures, empirically evaluating their effectiveness. Our comprehensive analysis identifies RINS as a powerful approach, outperforming +55 other methods like RAO used in Mobile LLM (Liu et al.,, 2024) and latent recurrent thinking (Geiping et al.,, 2025), and scaling inference by increasing the sequence length.
@@
    unlike prior works, we control for training compute FLOPs in our comparisons.
@@
    unlike prior works, we study the effectiveness of recursive depth beyond language to multimodal systems that incorporate language in their processing, such as contrastive models. In particular, our SigLIP-RINS-B/16 outperforms the popular SigLIP-B/16 (Zhai et al.,, 2023) by a wide margin; e.g. improving 0-shot accuracy in ImageNet from 77.3% to 79.6% and CIFAR100 from 70.3% to 80.7%.
@@
    argue that the performance gain of RINS likely stems from the self-similar (fractal) nature of language, by showing that a similar analysis in vision yields minimal improvements.
  5. 5.
@@
  6. 6.
    show that stochastic RINS can enhance performance even further, such as in multimodal systems, while offering the option to revert to non-recursive inference at test time with minimal performance degradation. In particular, with lightweight (linear) adapters, stochastic RINS offer a no-regret strategy, suggesting its potential as a viable component of LLM pretraining.
  7. 7.
    analyze the impact of KV cache sharing to reduce memory footprint even further and show that RINS continues to offer an advantage in that setup.
@@ 2 Recursive Inference Scaling
Overview.
Before describing how RINS works, we formalize definitions. Let $\mathcal{X}$ be a fixed domain, often the space of sequences of soft tokens of embedding dimension $d$. Let $\mathbb{L}=\{l_{1},l_{2},...,l_{n}\}$ be a fixed set of $n$ unique blocks, where each block $l_{k}:\mathcal{X}\to\mathcal{X}$ is a function mapping from the input space $\mathcal{X}$ to the same output space $\mathcal{X}$. By “unique” here we simply imply that such blocks (which typically comprise of multiple layers each) are not constrained to share the same parameters. Let $\mathcal{G}(\mathbb{L})$ be the space of all possible computation graphs representing neural network architectures that can be constructed by composing blocks from the set $\mathbb{L}$, while $f\in\mathcal{G}(\mathbb{L})$ be one specific architecture.
Figure 1 (right) illustrates some examples for the case when $|\mathbb{L}|=2$. For instance, one can repeatedly apply the entire model, as in the “repeat-all-over” (RAO) approach in Mobile LLM (Liu et al.,, 2024), or recursively apply a strict subset of the architecture, such as a single a block within the model. The choice of arrangement of blocks can significantly impact the model’s performance and efficiency.
Formally, let $C(f)$ be the actual computational cost (in FLOPs) of training $f\in\mathcal{G}(\mathbb{L})$ on a dataset sampled i.i.d. from distribution $\mathcal{D}$, considering only the forward pass. Also, $\mathcal{L}(f)$ is a performance metric of interest (e.g., validation loss) for model $f$, with lower values being better.
@@
For a fixed set of blocks $\mathbb{L}$ and a training compute budget $c$, a recursive architecture $f^{\star}\in\mathcal{G}(\mathbb{L})$ is called “better” than another $f\in\mathcal{G}(\mathbb{L})$ if $C(f^{\star})\leq C(f)\leq c$ and $\mathbb{E}[\mathcal{L}(f^{\star})]\leq\mathbb{E}[\mathcal{L}(f)]$.
@@
Model recursion offers a simple, plug-in approach for scaling inference time. By applying some layers iteratively to their own output, we effectively increase the computational path length during inference without altering the underlying model architecture. This allows us to exploit the benefits of increased inference compute. Importantly, it is complementary to other techniques like chain-of-thought (CoT) prompting (Wei et al.,, 2024) and repeated sampling (Li et al.,, 2022; Chen et al.,, 2021).
Figure 2: Language models are trained on 200B tokens. The $x$-axis is the training cost in units of layer $\times$ step. Notably, the performance advantage of RINS increases with longer training. The long-sequence baseline, using a context length of 1,536 tokens, exhibits lower performance due to processing fewer examples to maintain the same FLOPs count. See Figure 3(a) for longer training durations and Figure 4 for larger (1B parameter) models, further demonstrating the value of RINS. Sharp drops in perplexity near the end of training are due to learning rate cooldown.
@@
As discussed in Section 1, language has been shown to exhibit self-similarity, meaning that similar patterns repeat across multiple levels of granularity, from the structure of individual sentences to the composition of entire texts (Alabdulmohsin et al., 2024a, ). This observation suggests that recursive (scale-invariant) decoding could be particularly beneficial for language processing.
To determine the extent to which this is true, we systematically explore a vast space of parameter sharing strategies. We use the following taxonomy based on two key concepts: (1) signature and (2) degree. The “signature” of an architecture describes how parameter sharing is applied across different blocks of the model. For example, an architecture with the signature AB2C indicates that the model is partitioned into three unique blocks (A, B, and C) of equal size. The processing flow would be: apply block A on the input, apply B to A’s output, apply B again to its own output (hence the exponent), and apply C to the output of B. Note that while the signature AB2C has the same parameter count as ABC, it incurs $\approx 33\%$ more compute due to the repeated use of block B. This highlights the trade-off between computational cost and performance using recursion that we will carefully study in this work.
On the other hand, “degree” specifies if recursive depth is nested, as illustrated in Figure 1 (bottom). A degree of 2, for example, means that each of the blocks A, B, and C in our example have themselves the same recursive pattern used in the overall model. This adds another dimension to our taxonomy, allowing for a finer-grained classification of recursive architectures. Degree makes notation concise, but is not necessary; e.g. (ABB)2 is equivalent to ABBCDDCDD. By systematically varying both signature and degree, we can comprehensively explore the space of recursive models and identify optimal configurations. Appendix A provides a detailed pseudocode. With this framework, we now state our main question: Which family of architectures (i.e. signatures and degrees) lead to better performance according to Definition 2.1 under fixed compute budget?
To reiterate, this is a non-trivial question because it is possible for a non-recursive model to outperform all others given that it sees more training examples, since we always match from FLOPs. For instance, increasing the context length can be inferior to longer training within a compute constraint, as shown in Section 3. Nevertheless, our analysis reveals that Recursive INference Scaling (RINS) emerges as a clear winner. Its consistent superiority suggests that it captures a fundamental principle for efficiently processing language. We hypothesize that this is due to the self-similar geometry of language.
@@
RINS corresponds to architectures with degree 1 and signature ArB, for some $r>1$.
In other words, RINS partitions a model depthwise into two equally-sized blocks A and B. The first block A is recursively applied on its own output $r$ times before applying B. See Figure 1 for illustration and Appendix A for a detailed pseudocode. In Section 6, we study the optimal value of $r$.
@@
Our claim can be summarized as follows. Once a language model is chosen and the training compute budget is planned, one should enable RINS during training, which does not change the model size, and train the new RINS-enabled model for the same amount of compute. Our empirical results demonstrate that RINS-enabled models will consistently outperform baseline models. In addition, stochastic RINS, particularly with linear adapters (see Section 4), offers the option of forgoing scaling inference at test time with little performance degradation.